OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1).

OpenStudy (campbell_st):

find the gradient of the line the formula the formula for the gradient is \[m = \frac{(y _{2} - y _{1})}{(x _{2} -x _{1})}\] m = (-6 -1)/(-6 +3) \[m = \frac{-7}{-3} =========> m = \frac{7}{3}\] next the equation of the line in point slope form is y = mx + b or \[y = \frac{7x}{3} + b\] choice either point say ( -3, 1) substitute into the equation above and find b then you'll have the equation of the line